# GASLITEing the Retrieval: Exploring Vulnerabilities in Dense Embedding-based Search

### v1

@@ Appendix I On Practical Considerations in Poisoning a corpus
Finally, we note that even when the crafted input tokens are considered reversible, there could be more aspects relating the configuration of the tokenizer’s encoding that should be considered in a successful attack. For example, a tokenizer’s encoding may escape default tokenization of special tokens (e.g., to avoid encoding user’s text such as [PAD] as a padding token, but rather as the string ”[PAD]”).222222https://github.com/huggingface/transformers/pull/25081 232323https://x.com/karpathy/status/1823418177197646104 In this example, the performance of an attack that produces a passage with special tokens is expected to deteriorate. Notably, we find Cor.Pois. (Zhong et al., 2023) tend to create such samples (e.g., Tab. 9), however, we did not apply such escaping during evaluation. In response to these kind of issues, in our method, we do not allow the use of special tokens and non-printable tokens for crafting the passage. Our results (§6) show that GASLITE-like attacks are possible even if the recommended security practices, such as special-tokens escaping, are applied.

### v2

@@ Appendix H On Practical Considerations in Poisoning a corpus
First, poisoning a retrieval corpus is mostly done via insertion of the malicious text (e.g., uploading code section to a public repository or paragraphs to Wikipedia), as a result the input tokens depend on the tokenization process of the embedding model. Perhaps surprisingly, this capability is strictly weaker than directly controlling the input tokens, in particular, for many tokenizers (e.g., BERT’s; (Devlin et al., 2019)), there exist token lists that cannot be reached from any text, as demonstrated in Tab. 21. 292929https://gist.github.com/matanbt/afdd59f66158dc30b33d496864126e96
@@
Finally, we note that even when the crafted input tokens are considered reversible, there could be more aspects relating the configuration of the tokenizer’s encoding that should be considered in a successful attack. For example, a tokenizer’s encoding may escape default tokenization of special tokens (e.g., to avoid encoding user’s text such as [PAD] as a padding token, but rather as the string ”[PAD]”).303030https://github.com/huggingface/transformers/pull/25081313131https://x.com/karpathy/status/1823418177197646104 In this example, the performance of an attack that produces a passage with special tokens is expected to deteriorate. Notably, we find Cor.Pois. (Zhong et al., 2023) tend to create such samples (e.g., Tab. 15), however, we did not apply such escaping during evaluation. In response to these kind of issues, in our method, we do not allow the use of special tokens and non-printable tokens for crafting the passage. Our results (§6) show that GASLITE-like attacks are possible even if the recommended security practices, such as special-tokens escaping, are applied.